OpenStudy (anonymous):

Write down an equation: The product of 2 and 3 more than a number is 7.

OpenStudy (anonymous):

x+3 = (3 more than a number)

OpenStudy (anonymous):

2(x+3) = (The product of 2 and 3 more than a number)

OpenStudy (anonymous):

so the equation is 2(x+3)=7

OpenStudy (anonymous):

thankyou